1.   What will happen to the equilibrium of tomatoes if the number of sellers increase in the market?



1
Expert's answer
2021-02-18T07:54:33-0500

The number of sellers impacts on the quantity of tomatoes that can be supplied during some period. It is well known that the firm's supply function shows the quantity to be produced with respect to the different prices (during some period). Let's assume that in initial state (denoted by S1) N1 firms provide tomatoes. In the second state (denoted by S2) N2 firms provide tomatoes (for example, N2>N1). Let's assume that each firm provide the tomatoes according to one supply function Qs(P). Hence, the market supply function in the state S1 can be written as follows "MQs(P)=N1\times Qs(P)"

In the state S2 the supply function can be written as follows:

"MQs(P)=N2\times Qs(P)".

Let market demand on tomatoes be denoted by MD.

The equilibrium level in ste state 1 can be calculated acvording to the equation:

"MD(P)=MQs(P)=N1\times Qs(P)".

In the state 2:

"MD(P)=MQs(P)=N2\times Qs(P)".

Hence, the equilibrium level of Q will be increased:

"MD(P)\/N1>MD(P)\/N2" .

In the axis P and Q (Q is vertical) the supply curve will be shifted to left. If the demand curved is not shifted, the equilibrium level of Q will be increased. The equilibrium level of price will be decreased (according to the market demand curve).

ANSWER:

The equilibruim level of tomatoes sold will be increased. The equilibrium level of price will be decreased.